OpenStudy (anonymous):

A magic show sold tickets for $2 for every child and $3 for every adult. The total sales were $118. The difference of two times the number of adults and half the children will total 20. Using the elimination method for the system of equations, what is the number of adults and children at the magic show? A. adults = 5 children = 30 B. adults = 18 children = 32 C. adults = 16 children = 35 D. adults = 32 children = 18

OpenStudy (anonymous):

Are you having trouble setting up the equations?

OpenStudy (anonymous):

I am working on very simliar equations..I know how to set up the equations but dont know how to work them once I have them...

OpenStudy (anonymous):

I know I need to eliminate a variable...but not sure how..I have 40x + 100y=125 and 80x + 84y= 163.

OpenStudy (anonymous):

@NCMom post it as your own question and I or someone else will help

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@CalebBeavers yes!

OpenStudy (anonymous):

c=child tickets a=adult tickets 2c+3a=118 2a-1/2c=20 can you solve now?

OpenStudy (anonymous):

So i have to divide @CalebBeavers ?

OpenStudy (anonymous):

No...use elimination or substitution to solve the system

OpenStudy (anonymous):

@CalebBeavers Sorry, math just isnt my subject lol im actually kinda confused!

OpenStudy (anonymous):

Solve for a in one of the equations

OpenStudy (anonymous):

@CalebBeavers how do i use the elimination system in this?